OpenStudy (anonymous):

a long straight rope is lying on a floor is pulled along its length with uniform acceleration along its lenght if the velocities of the front and back ends of the rope when it passes through a particular point on the floor are 2m/s and 4m/s respectively the velocity of the midpoint of the rope when it passes through the same point is?

OpenStudy (experimentx):

\( v_1 = 4, u = 2 \) \( (v_1)^2 = u^2 + 2as \) \( (v_2)^2 = u^2 + 2a(s/2) \) Find \( v_2 \) !!!

OpenStudy (anonymous):

did nt understand (v2)2=u2+2a(s/2)

OpenStudy (experimentx):

the distance travelled by midpoint will be half the distance travelled by the rear end.

OpenStudy (anonymous):

ok now i get it thanxxx